# Online Episodic Memory Visual Query Localization with Egocentric Streaming Object Memory

### v1

@@ 6.2 Analysis of ESOM and Ablations
[…]
@@
Figure 6: Qualitative example. Successful retrieval of query $\mathbf{Q}$ from $\mathcal{M}_{ego}$ is shown using (a) the oracle OMP algorithm (ODoracle - OToracle) and (b) the real-world OMP (ODyolo - OTegostark). The figure highlights the visual representations stored in $\mathcal{M}_{ego}$ and matched during retrieval, their respective timestamps, and the duration of the response track $\mathbf{r}$.
@@ 7 Conclusions
We introduced the novel task of Online Episodic Memory - Visual Query Localization (OEM-VQL), where an object query must be localized in the past without access to the original video. To address this task, we proposed ESOM (Egocentric Stream Object Memory), a new online approach that memorizes only essential objects to support efficient visual query retrieval and localization. ESOM significantly reduces memory and computational demands, achieving up to a 1000$\times$ reduction in memory usage and retrieval times from minutes to seconds. Experiments on the Ego4D dataset show ESOM’s superior accuracy (81.92% with oracular models) over offline methods (55.89%), though real-world performance is still limited by current detection and tracking technology. Besides establishing a benchmark for OEM-VQL, we also contribute a principled evaluation, based on a real downstream task, for object detection and tracking algorithms for egovision applications.
Acknowledgements.
This work was funded by European Union- NextGenerationEU under Project PRIN 2022 PNRR TEAM.
[…]
@@ Acknowledgements.
This document reports implementation details on the proposed ESOM methodology, along with additional results and discussion. Specifically, Section A explains the ESOM architecture by providing details of the Object Memory representation (sub-section A.1), the Object Tracking (OT) module (sub-section A.2.2), on the Object Discovery (OD) module (Section A.2.2), and on on the Query Retrieval and Localization (QRL) algorithm (sub-section A.3). Section  B reports details on the implementation of the baselines methods on AMEGO [17] and VideoAgent [11]. Section C reports information on the evaluation dataset and Section D reports additional and qualitative results.
@@ A.1 Object Memory
Memory representation.
We defined the object memory $\mathcal{M}_{ego}$ as a set of of high-level representations $\mathbf{O}_{i}$ that capture the spatio-temporal locations of the objects in the frames $\mathbf{F}_{t}$ of a video $\mathcal{V}_{ego}$. Each index $1\leq i\leq N$ corresponds to a unique object instance. The object representation $\mathbf{O}_{i}$ is structured as a list of records, temporally indexed by $t$, where each record having the following form: $\mathbf{o}_{i,t}=[\mathbf{b}_{i,t},\phi_{i,t}]$. Here $\mathbf{b}_{i,t}=[b_{i,t}^{x},b_{i,t}^{y},b_{i,t}^{w},b_{i,t}^{h}]$ denotes the object’s bounding box, and $\phi_{i,t}$ represents the object’s visual representation at time $t$ as the RGB patch outlined by the bounding-box $\mathbf{b}_{i,t}$. The design of $\mathcal{M}_{ego}$ was driven by experimental results. Indeed, we evaluated various memory representations, focusing on four distinct forms of objects representation $\mathbf{O}_{i}$:
  • MR1* (design as reported in Section 4.1 of the main paper): for every object $\mathbf{O}_{i}$, the memory $\mathcal{M}_{ego}$ retains the RGB patch $(\phi_{i,t})$ and $\mathbf{b}_{i,t}$ of the object at the moment of object discovery, as well as two RGB patches belonging to the most recent response track. These patches are chosen as the first and last valid patches identified by the classifier within the most recent sequence of bounding-boxes. For all other temporal locations not mentioned above, the memory $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$.
  • MR2: for every object $\mathbf{O}_{i}$, the memory $\mathcal{M}_{ego}$ retains the RGB patch $(\phi_{i,t})$ and $\mathbf{b}_{i,t}$ of the object at the moment of object discovery, as well as all the RGB patches $(\phi_{i,t})$ and bounding-boxes $\mathbf{b}_{i,t}$ belonging to the most recent response track, considered useful by the classifier. For all other frames in temporal locations not mentioned, the memory $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$. This implementation was conducted to assess the impact of exploiting more object visual information for QRL.
  • MR3: for every object $\mathbf{O}_{i}$, the memory $\mathcal{M}_{ego}$ retains the RGB patch $(\phi_{i,t})$ and $\mathbf{b}_{i,t}$ of the object at the moment of object discovery. For all the other temporal locations, the memory $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$. This implementation was conducted to assess the impact of exploiting less object visual information for QRL, thus to reduce the size and processing time of $\mathcal{M}_{ego}$.
  • MR4: for every object $\mathbf{O}_{i}$, the memory $\mathcal{M}_{ego}$ retains the RGB patch $(\phi_{i,t})$ and $\mathbf{b}_{i,t}$ of the object at the moment of object discovery as well as the RGB patches $(\phi_{i,t})$ and $\mathbf{b}_{i,t}$ in the first set of contiguous bounding-boxes (i.e., the beginning of each tracklet). For all the other temporal locations, the memory $\mathcal{M}_{ego}$ stores only the spatio-temporal information provided by $\mathbf{b}_{i,t}$. This implementation was conducted to assess the impact of exploiting object visual information for QRL that refers to a better performance of the OT module.
The OEM-VQL performance of these memory configurations built by the Object Memory Population (OMP) algorithm and processed by the Query Retrieval and Localization (QRL) algorithm are reported in Table 6. As can be noticed, the proposed memory representation (MR1*) offers the best balance between QRL accuracy and memory size and retrieval time.
@@
To create a consistent yet compact memory, the .write() operation incorporates a filtering mechanism to determine which pieces of object information are stored in $\mathcal{M}_{ego}$. This filtering process relies on a binary classification function that assigns a probability score to each visual representation $\phi_{i,t}$, assessing its quality for effective query retrieval. The classification model was implemented using a deep neural network consisting of a ResNet18 [19] feature extractor and a linear layer, which served as the representation filter. This model was fine-tuned on a dataset of 17071 visual representations extracted from 250 uniformly sampled clips from the training set of the Ego4D Episodic Memory VQ2D-EgoTracks benchmark [18, 42]. Specifically, the training set included 8534 visual representations labeled as 0 (i.e., not suitable for QRL) and 5153 representations labeled as 1 (i.e., suitable for QRL). The validation set consisted of 2126 visual representations labeled as 0 and 1258 representations labeled as 1.
@@
The training and validation representations were labeled as 0 or 1 based on whether they were incorrectly or correctly matched by the QRL algorithm. For each frame in the 250 selected clips, the representation $\phi_{i}$ of the annotated visual query $\mathbf{Q}_{i}$ was extracted, and its feature representation $\Phi(\phi_{i})$ was computed using DINOv2 [30] as $\Phi(\cdot)$. This representation was then compared to $\Phi(\phi_{i,t})$, the feature representation of the patches withing the ground-truth bounding-boxes of the tracklet associated with the query $\mathbf{Q}_{i}$, using cosine similarity. This comparison produced a similarity score $s_{t}$ ranging from 0 to 1. Patches were assigned negative labels if $s_{t}\leq 0.3$ and positive labels if $s_{t}\geq 0.7$. After labeling, the data was randomly split into training and validation sets, with 80% allocated to the training set and 20% to the validation set.
@@
The visual representation classifier was fine-tuned for 50 epochs using Stochastic Gradient Descent (SGD) with a learning rate of 0.01. At convergence, it achieved a test accuracy of  96%. The input patches were resized to $224\times 224$ pixels before being fed into the model.
OMP Configuration Memory tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$
@@
MR1* 0.73 0.68 81.92 7.1 MB
MR2 0.73 0.68 81.92 395 MB
MR3 0.63 0.55 74.33 2.6 MB
ODoracle OToracle MR4 0.79 0.71 83.38 38.5 MB
MR1* 0.21 0.19 40.55 11.2 MB
MR2 0.21 0.19 40.55 94.0 MB
MR3 0.20 0.17 38.02 1.0 MB
ODyolo OToracle MR4 0.22 0.19 41.82 100.8 MB
MR1* 0.18 0.05 31.91 40.3 MB
MR2 0.18 0.05 31.91 350.0 MB
MR3 0.16 0.03 26.56 3.1 MB
ODoracle OTegostark MR4 0.16 0.04 29.02 4.1 GB
MR1* 0.0003 0.002 4.02 209.3 MB
MR2 0.0003 0.002 4.02 2.3 GB
MR3 0.0002 0.0003 1.78 105.1 MB
ODyolo OTegostark MR4 0.0002 0.0005 2.90 746.6 MB
Table 6: Effect of memory representation in ESOM. This table illustrates the impact of using different representations for $\mathcal{M}_{ego}$ on QRL performance and memory size. MR1* refers to the representation described in the main paper. As can be noticed, MR1* constitutes the best approach at the intersection of QRL accuracy and memory efficiency. (Best results are highlighted in bold).
@@
MASA GroundingDino 0.036 0.211 0.105 0.766
MASA YOLOv10* 0.066 0.080 0.116 0.798
@@
EgoSTARK Hand-Object Detector 0.034 0.121 0.102 0.739
EgoSTARK GroundingDino 0.057 0.121 0.115 0.803
EgoSTARK YOLOv10* 0.061 0.188 0.151 0.801
EgoSTARK Ground-truth 0.217 0.390 0.347 0.831
Ground-truth YOLOv10* 0.848 0.848 0.943 1.000
Ground-truth GroundingDino 0.875 0.875 0.943 1.000
@@ A.2 Object Memory Population (OMP)
OTegostark implements a multiple-object tracking approach using multiple instances of single-object trackers. This approach is motivated by the lack of effective multiple-object tracking solutions for egocentric vision, whereas single-object tracking has been extensively studied [9, 10, 42]. For single-object tracking, we use the STARK framework [48], fine-tuned on the EgoTracks dataset as described in [42], resulting in the EgoSTARK tracker. To track multiple objects, we instantiate multiple EgoSTARK trackers simultaneously. A new tracker instance is initialized in frame $\mathbf{F}_{t}$ whenever the OD module identifies a valid object, which is then stored in memory. Each bounding-box predicted by the OD module serves as a template to initialize a corresponding tracker. Once initialized, a tracker runs across subsequent frames to predict the bounding boxes of the object it was assigned to, thereby updating the spatiotemporal representation of the respective object. Due to the challenges of object tracking in egocentric vision [9, 10, 42], we found different tracker instances to often end in tracking the same object. In order to prevent duplicate tracker instances, we calculate the IoU between all tracked objects at frame $t$. If two tracking predictions overlap with an IoU greater than $0.5$, the trackers are considered to be duplicate trackers and only the tracker instance and their corresponding prediction with the lower object instance ID (i.e. the older object) is kept.
Furthermore, to prevent incorrect tracking predictions from causing a tracker to deviate from a searching ares that contains the target object and wander across the frame, the box used to compute the search region for the subsequent frame is not updated if the confidence score of the EgoSTARK instance falls below 0.4. This strategy is based on the observation that, in egocentric vision, objects often leave the frame during rapid head movements but typically reappear in a similar location when the head returns to its original position. Before writing the OT module predictions to $\mathcal{M}_{ego}$, the bounding-boxes $\mathbf{b}_{i,t}$ with confidence score $s_{i,t}>\lambda_{ot}=0.5$ are retained.
@@ A.2.2 Object Discovery (OD) module
As detecting objects from egocentric videos is not straightforward [51], we tested different object detectors including the off-the-shelf detectors GroundingDino [26] and the hand-object detector of [39] (here referred to as Hand-Object Detector) which do not require training on the considered settings, as well as Faster R-CNN R101-FPN [36] and YOLOv10 [45] instances trained on different data sources, including EgoObjects [51] and curated versions of EgoTracks [42]. Targeting visual object tracking, EgoTracks contains spatially-sparse bounding box annotations across contiguous frames (i.e., only few object instances are annotated in each frame). Each object instance is associated to a textual description rather than an object class (e.g., “a blue bottle”). To enable the training of object detectors, we first clustered object descriptions with K-Means to obtain an initial set of $1169$ classes, followed by manual refinement in order to adjust them, leading to a diverse, granular taxonomy comprising $295$ classes.
Figure 7: Visualization of discarded bounding-boxes. Small-area bounding-boxes similar to those shown in this figure have been discarded for the training and evaluation of the module. This is because we found that they lacked sufficient visual information, leading to an increased number of false positive detections. These detections provided no useful data for tracker initialization within the OT module or for effective QRL.
@@
Since EgoTracks contains annotations in contiguous frames, we uniformly sampled one third of all frames. EgoTracks contains annotations for very small objects to assess the ability of visual object trackers to continuously follow objects subject to scale changes. Since these objects would be unfeasible to detect from single frames and may be source of ambiguity due to occlusion or motion blur, we discarded bounding box annotations with areas falling in the first quartile of all bounding box areas in the training set. Figure 7 shows a few examples of annotation bounding boxes that fall in the first quartile and therefore discarded from the dataset. This leads to a reduced training set of $474,761$ annotated images, while we consider a validation set of $2,466$ images.
@@
YOLOv10 - based detector.
A second version of the object discovery module was based on the YOLOv10 [45] object detector, given its capacity for real-time inference. We trained the YOLO detector on the same set of refined EgoTracks annotations. Given the sparsity of EgoTracks annotations, both the Faster R-CNN and YOLO detectors tend to recognize few object instances per image, which may limit the ability of the object discovery module to identify all possible objects of interest. To address this issue, we trained the YOLO object detector on the EgoObjects [51] dataset, which contains dense object annotations for egocentric frames with a fixed taxonomy of $638$ classes. To make the EgoObjects and EgoTracks sets compatible, we mapped each EgoTracks object description to its corresponding class from the EgoObjects taxonomy. This is done by first describing the object description and class name with a Sentence Transformer, and then selecting the class maximizing the cosine similarity between the embeddings. To reduce noise, we discarded all annotations with a similarity lower than $0.7$.
@@
Table 8 shows some examples of object descriptions retrieved from the EgoTracks annotations and mapped to EgoObjects objects. Note that some objects such as “air conditioner”, “headphones” or “yoga mat” have a reasonable match, while “washing machine” contains “flat washer”, which has a different meaning despite a close embedding. We did not manually adjust these inconsistencies to keep the process fully automated. The table contains one class with no matches (accordion), due to the absence of EgoTracks objects with similar descriptions. Out of the total number of class descriptions (3491), we mapped 1241 captions, while the remaining 2250 classes were excluded from the dataset. This resulted in a high-quality subset of the data suitable for training and evaluation.
@@
Despite the domain difference between EgoObjects and EgoTracks, we noticed improved qualitative and quantitative results. To further adapt the model to the EgoTracks domain, we fine-tuned the YOLO object detector, pre-trained on EgoObjects, on the EgoTracks dataset with labels mapped to the EgoObjects taxonomy. Table 9 compares the performance of the considered approaches with two off-the-shelf object detectors: GroundingDino [26] and the Hand-Object-Detector of [39]. We do not train these detectors on the EgoTracks data and use the official checkpoints and implementations provided by the authors. Since Visual Query Localization does not require to correctly identify object classes, all methods are evaluated with AP and Recall, merging all ground truth and predicted classes in a single “object” class. It is worth noting that AP and Recall measure different aspects of the detector. Indeed, AP measures the average ratio between correct predictions and all predictions made by the detector, whereas Recall measures the fraction of objects which have been correctly found by the detector, regardless of the total number of objects. In our settings, it is desirable to have a large recall (future queries have more chance to end up in the memory), but, since it is not possible to track all objects and a select must be made based on the confidence score, in practice, a very low value of AP is not desirable. The harmonic mean between AP and Recall is hence adopted as a way to identify the most balanced detector. Results in Table 9, obtained with a confidence score threshold $\lambda_{od}=0.01$, show that Faster R-CNN trained on labels clustered from EgoTracks achieves the highest recall ($0.5224$), but also yields a very low AP value ($0.0219$). On the contrary, the YOLO instance achieves the highest AP ($0.1983$) but also a low recall ($0.1634$). We attribute these skewed results to the sparse nature of annotations in EgoTracks (only few object instances annotated per frame) and the different ways of the two methods to handle negative examples. The harmonic mean results of the first one are rather low ($0.0420$), while the second result is higher but not optimal at $0.1791$, highlighting the unbalanced nature of the two approaches. Training YOLO on EgoObjects, which contains much denser annotations increases Recall to $0.3780$ at the expense of AP ($0.0247$), which leads to a low harmonic mean of $0.0463$. Mapping EgoTracks annotations to the EgoObjects taxonomy and training YOLO using both datasets provides the most balanced detector, with an harmonic mean of $0.1882$, an AP of $0.1358$ and a Recall of $0.3067$. GroundingDino achieves a low AP and a reasonable recall value ($0.0450$ and $0.2572$), which is likely due to the large domain shift between data used for training GroundingDino and the challenging egocentric observations of EgoTracks. Using Hand-Object-Detector as an object detector leads to lower results, which is due to the fact that the Hand-Object-Detector only focuses on interacted objects, hence neglecting many non-interacted objects which may nevertheless be queried in the future. Lastly, RT-DETR (used in VideoAgent [11]) achieves results comparable to GroundingDino, due to similar reasons.

### v2

@@ 6.3 Detailed Analysis of ESOM and Ablations
Figure 5 shows retrieval performance, storage space, and retrieval time as 25%, 50%, 75%, and 100% of videos are observed with different OMP configurations. QRL performance declines slightly with video length and remains steady with oracular OD and OT, while minor drops occur with ODoracle and ODegostark, underscoring ESOM’s scalability. Storage and retrieval time generally increase over time; however, memory size remains stable with ODoracle since most objects appear within the first 25% of the video. ODyolo - OTegostark significantly impacts storage and retrieval due to duplicate objects from weaker tracking.
@@
Figure 6: Qualitative example. Successful retrieval of query $\mathbf{Q}$ from $\mathcal{M}_{ego}$ using the oracle OMP algorithm (ODoracle - OToracle) and the real-world OMP (ODyolo - OTegostark). Sample frames from each track are shown with corresponding bounding boxes whose patches $\phi_{i,t}$ were matched during retrieval. The frame $t=0$ corresponds to the moment of the object’s discovery. The green bar visualizes the temporal extent of the ground-truth track, with timestamps indicating its duration.
@@ 7 Conclusions
In the paper, we introduced the novel task of Online Visual Queries 2D (OVQ2D), where a video is processed in a streaming fashion and an object query must be localized without access to the original video. To address this problem, we proposed ESOM (Egocentric Stream Object Memory), a new online approach that detects, tracks, and memorizes essential objects to support efficient visual query retrieval and localization. ESOM significantly optimizes both memory and computational efficiency, reducing memory usage by up to 7$\times$. Additionally, it shortens retrieval times, bringing them down from minutes or tens of seconds to just a few seconds. Experiments on the Ego4D dataset show ESOM’s superior performance when compared with other online methods and its potential when object detection and tracking components are improved. Besides establishing a benchmark for OVQ2D, we also contribute a principled evaluation, based on a real downstream task, for object detection and tracking algorithms for egovision applications.
Acknowledgments.
Progetto PRIN 2022 PNRR - “Tracking in Egovision for Applied Memory (TEAM)” Codice P20225MSER_001. Codici CUP G53D23006680001 (Università di Udine) e E53D23016240001 (Università di Catania). This research has been funded by the European Union, NextGenerationEU – PNRR M4 C2 I1.1 RS Micheloni.
@@ Acknowledgments.
This document reports implementation details on the proposed ESOM methodology, along with additional results and discussion. Specifically, Section A explains the ESOM architecture by providing details of the Object Memory representation (sub-section A.1), the Object Tracking (OT) module (sub-section A.2.2), on the Object Discovery (OD) module (Section A.2.2), and on the Query Retrieval and Localization (QRL) algorithm (sub-section A.3). Section  B reports details on the implementation of the baselines methods on AMEGO [21] and VideoAgent [15]. Section C reports information on the evaluation dataset and Section D reports additional and qualitative results.
@@ A.1 Object Memory
To improve retrieval time and accuracy, the .write() operation incorporates a filtering mechanism to determine whether the visual content within $\mathbf{b}_{i,t}$, $\phi_{i,t}$, is relevant for the retrieval process. This filtering process relies on a binary classification function that assigns a label $c_{i,t}$ to each visual representation $\phi_{i,t}$, assessing its quality for effective query retrieval.
The design of $\mathcal{M}_{ego}$, which includes information about the quality of data for retrieval, was guided by experimental results. Specifically, we tested various filtering approaches to determine their effectiveness.
  • MR1* (design as reported in Section 4.1 of the main paper): for each object $\mathbf{O}_{i}$, the filter assigns $c_{i,t}=1$ at the moment of object discovery, as well as to two bounding-boxes $\mathbf{b}_{i,t}$ and their associated frames $\mathbf{F}_{t}$ from the most recent response track. These bounding-box and frame pairs are selected as the first and last valid patches identified by the classifier within the latest sequence of bounding boxes. For all other frames in temporal locations not mentioned, the filter set $c_{i,t}=0$.
  • MR2: for each object $\mathbf{O}_{i}$, the filter assigns $c_{i,t}=1$ at the moment of object discovery, as well as to all the bounding-boxes $\mathbf{b}_{i,t}$ and their associated frames $\mathbf{F}_{t}$ from the most recent response track, considered useful by the classifier. This implementation was conducted to assess the impact of exploiting more object visual information for QRL.
  • MR3: for each object $\mathbf{O}_{i}$, the filter assigns $c_{i,t}=1$ at the moment of object discovery. For all other frames in temporal locations not mentioned, the filter set $c_{i,t}=0$. This implementation was conducted to assess the impact of exploiting less object visual information for QRL, thus reducing the retrieval time of $\mathcal{M}_{ego}$.
  • MR4: for each object $\mathbf{O}_{i}$, the filter assigns $c_{i,t}=1$ to two bounding-boxes $\mathbf{b}_{i,t}$ and their associated frames $\mathbf{F}_{t}$ from the most recent response track. These bounding-box and frame pairs are selected as the first and last valid patches identified by the classifier within the latest sequence of bounding boxes. For all other frames in temporal locations not mentioned, the filter set $c_{i,t}=0$. This implementation was conducted to assess the impact of considering the object at the moment of discovery in retrieval performances.
The OVQ2D performance of these memory configurations built by the Object Memory Population (OMP) algorithm and processed by the Query Retrieval and Localization (QRL) algorithm are reported in Table 6. As can be noticed, the proposed memory representation (MR1*) offers the best balance between QRL accuracy and retrieval time.
@@
Classifier implementation details The classification model was implemented using a deep neural network consisting of a ResNet18 [23] feature extractor and a linear layer, which served as the representation filter. This model was fine-tuned on a dataset of 17,071 visual representations extracted from 250 uniformly sampled clips from the training set of the Ego4D Episodic Memory VQ2D-EgoTracks benchmark [22, 46]. Specifically, the training set included 8,534 visual representations labeled as 0 (i.e., not suitable for QRL) and 5,153 representations labeled as 1 (i.e., suitable for QRL). The validation set consisted of 2,126 visual representations labeled as 0 and 1,258 representations labeled as 1.
@@
The training and validation representations were labeled as 0 or 1 based on whether they were incorrectly or correctly matched by the QRL algorithm. For each frame in the 250 selected clips, the representation $\phi_{i}$ of the annotated visual query $\mathbf{Q}_{i}$ was extracted, and its feature representation $\Phi(\phi_{i})$ was computed using Siam-RCNN [48] as $\Phi(\cdot)$. This representation was then compared to $\Phi(\phi_{i,t})$, the feature representation of the patches withing the ground-truth bounding boxes of the tracklet associated with the query $\mathbf{Q}_{i}$, using a classification head [48]. This comparison produced a similarity score $s_{t}$ ranging from 0 to 1. Patches were assigned negative labels if $s_{t}\leq 0.3$ and positive labels if $s_{t}\geq 0.7$. After labeling, the data was randomly split into training and validation sets, with 80% allocated to the training set and 20% to the validation set.
@@
The visual representation classifier was fine-tuned for 50 epochs using Stochastic Gradient Descent (SGD) with a learning rate of 0.01. At convergence, it achieved a test accuracy of  97%. The input patches were resized to $224\times 224$ pixels before being fed into the model.
OMP Configuration Memory %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Time $\downarrow$
@@
MR1* 73.2 68.1 81.92 0.51 s
MR2 73.3 68.3 81.92 4.27 s
MR3 63.7 55.1 74.33 0.05 s
ODoracle OToracle MR4 70.4 66.2 80.35 0.32 s
MR1* 21.0 19.4 40.55 0.36 s
MR2 21.0 19.5 40.55 2.71 s
MR3 20.3 17.0 38.02 0.05 s
ODyolo OToracle MR4 20.7 17.2 39.05 0.21 s
MR1* 0.18 0.05 31.91 0.49 s
MR2 0.18 0.05 31.91 4.98 s
MR3 0.16 0.03 26.56 0.05 s
ODoracle OTegostark MR4 0.16 0.04 29.02 0.29 s
MR1* 0.03 0.02 4.02 3.08 s
MR2 0.03 0.02 4.02 14.90 s
MR3 0.02 0.03 1.78 0.05 s
ODyolo OTegostark MR4 0.02 0.05 2.90 0.69 s
Table 6: Effect of different filtering strategies. This table illustrates the impact of using different filtering approaches for $\mathcal{M}_{ego}$ population on QRL performance and retrieval time. MR1* refers to the representation described in the main paper. As can be noticed, MR1* constitutes the best approach at the intersection of QRL accuracy and retrieval efficiency. (Best results are highlighted in bold).
@@
MASA GroundingDino 0.036 0.211 0.105 0.013
MASA YOLOv10* 0.066 0.080 0.116 0.038
@@
EgoSTARK Hand-Object Detector 0.034 0.121 0.102 0.012
EgoSTARK GroundingDino 0.057 0.121 0.115 0.029
EgoSTARK YOLOv10* 0.061 0.188 0.151 0.025
EgoSTARK Ground-truth 0.217 0.390 0.347 0.137
BoT-SORT YOLOWorld 0.014 0.073 0.042 0.005
Ground-truth YOLOv10* 0.848 0.848 0.943 0.762
Ground-truth GroundingDino 0.875 0.875 0.943 0.812
@@ A.2 Object Memory Population (OMP)
OTegostark implements a multiple-object tracking approach using multiple instances of single-object trackers. This approach is motivated by the lack of effective multiple-object tracking solutions for egocentric vision, whereas single-object tracking has been extensively studied [11, 12, 46]. For single-object tracking, we use the STARK framework [52], fine-tuned on the EgoTracks dataset as described in [46], resulting in the EgoSTARK tracker. To track multiple objects, we instantiate multiple EgoSTARK trackers simultaneously. A new tracker instance is initialized in frame $\mathbf{F}_{t}$ whenever the OD module identifies a valid object, which is then stored in memory. Each bounding box predicted by the OD module serves as a template to initialize a corresponding tracker. Once initialized, a tracker runs across subsequent frames to predict the bounding boxes of the object it was assigned to, thereby updating the spatio-temporal representation of the respective object. Due to the challenges of object tracking in egocentric vision [11, 12, 46], we found different tracker instances to often end in tracking the same object. In order to prevent duplicate tracker instances, we calculate the IoU between all tracked objects at frame $t$. If two tracking predictions overlap with an IoU greater than $0.5$, the trackers are considered to be duplicate and only the tracker instance and their corresponding prediction with the lower object instance ID (i.e. the older object) is retained.
Furthermore, to prevent incorrect tracking predictions from causing a tracker to deviate from a searching area that contains the target object and wander across the frame, the box used to compute the search region for the subsequent frame is not updated if the confidence score of the EgoSTARK instance falls below 0.4. This strategy is based on the observation that, in egocentric vision, objects often leave the frame during rapid head movements but typically reappear in a similar location when the head returns to its original position. Before writing the OT module predictions to $\mathcal{M}_{ego}$, the bounding boxes $\mathbf{b}_{i,t}$ with confidence score $s_{i,t}>\lambda_{ot}=0.5$ are retained.
@@
OTyoloworld.
Finally, we investigated the tracking algorithm BoT-SORT [1] included in the YOLOWorld [8] implementation.
OToracle
We implement the oracular object tracker as follows: whenever a detected object intersects a GT bounding box from EgoTracks, we put the rest of the track (from that box onward) in memory. EgoTracks extends the Episodic Memory VQ2D benchmark introducing an object track for each query, hence GT tracks from EgoTracks are valid GT VQ2D tracks. When we combine OToracle with ODoracle, we simply insert in memory all object tracks from EgoTracks.
@@ A.2.2 Object Discovery (OD) module
As detecting objects from egocentric videos is not straightforward [57], we tested different object detectors including the off-the-shelf detectors GroundingDino [30] and the hand-object detector of [43] (here referred to as Hand-Object Detector) which do not require training on the considered settings, as well as Faster R-CNN R101-FPN [40] and YOLOv10 [49] instances trained on different data sources, including EgoObjects [57] and curated versions of EgoTracks [46]. Targeting visual object tracking, EgoTracks contains spatially sparse bounding box annotations across contiguous frames (i.e., only few object instances are annotated in each frame). Each object instance is associated to a textual description rather than an object class (e.g., “a blue bottle”). To enable the training of object detectors, we first clustered object descriptions with K-Means to obtain an initial set of $1,169$ classes, followed by manual refinement in order to adjust them, leading to a diverse, granular taxonomy comprising $295$ classes.
Figure 7: Visualization of discarded bounding-boxes. Small-area bounding boxes similar to those shown in this figure have been discarded for the training and evaluation of the module. This is because we found that they lacked sufficient visual information, leading to an increased number of false positive detections. These detections provided no useful data for tracker initialization within the OT module or for effective QRL.
@@
Since EgoTracks contains annotations in contiguous frames, we uniformly sampled one-third of all frames. EgoTracks contains annotations for very small objects to assess the ability of visual object trackers to continuously follow objects subject to scale changes. Since these objects would be unfeasible to detect from single frames and may be source of ambiguity due to occlusion or motion blur, we discarded bounding box annotations with areas falling in the first quartile of all bounding box areas in the training set. Figure 7 shows a few examples of annotation bounding boxes that fall in the first quartile and therefore discarded from the dataset. This leads to a reduced training set of $474,761$ annotated images, while we consider a validation set of $2,466$ images.
@@
YOLO - based detector.
A second version of the object discovery module was based on the YOLOv10 [49] object detector, given its capacity for real-time inference. We trained the YOLOv10 detector on the same set of refined EgoTracks annotations. Given the sparsity of EgoTracks annotations, both the Faster R-CNN and YOLO detectors tend to recognize a few object instances per image, which may limit the ability of the object discovery module to identify all possible objects of interest. To address this issue, we trained the YOLOv10 object detector on the EgoObjects [57] dataset, which contains dense object annotations for egocentric frames with a fixed taxonomy of $638$ classes. To make the EgoObjects and EgoTracks sets compatible, we mapped each EgoTracks object description to its corresponding class from the EgoObjects taxonomy. This is done by first describing the object description and class name with a Sentence Transformer, and then selecting the class maximizing the cosine similarity between the embeddings. To reduce noise, we discarded all annotations with a similarity lower than $0.7$.
@@
Table 8 shows some examples of object descriptions retrieved from the EgoTracks annotations and mapped to EgoObjects objects. Note that some objects such as “air conditioner”, “headphones” or “yoga mat” have a reasonable match, while “washing machine” contains “flat washer”, which has a different meaning despite a close embedding. We did not manually adjust these inconsistencies to keep the process fully automated. The table contains one class with no matches (accordion), due to the absence of EgoTracks objects with similar descriptions. Out of the total number of class descriptions (3491), we mapped 1,241 captions, while the remaining 2,250 classes were excluded from the dataset. This resulted in a high-quality subset of the data suitable for training and evaluation.
@@
Despite the domain difference between EgoObjects and EgoTracks, we noticed improved qualitative and quantitative results. To further adapt the model to the EgoTracks domain, we fine-tuned the YOLOv10 object detector, pre-trained on EgoObjects, on the EgoTracks dataset with labels mapped to the EgoObjects taxonomy. Table 9 compares the performance of the considered approaches with three off-the-shelf object detectors: GroundingDino [30], the Hand-Object-Detector of [43], and YoloWorld [8]. We do not train these detectors on the EgoTracks data and use the official checkpoints and implementations provided by the authors. Since Visual Query Localization does not require to correctly identify object classes, all methods are evaluated with AP and Recall, merging all ground truth and predicted classes in a single “object” class. It is worth noting that AP and Recall measure different aspects of the detector. Indeed, AP measures the average ratio between correct predictions and all predictions made by the detector, whereas Recall measures the fraction of objects which have been correctly found by the detector, regardless of the total number of objects. In our settings, it is desirable to have a large recall (future queries have more chance to end up in the memory), but, since it is not possible to track all objects and a selection must be made based on the confidence score, in practice, a very low value of AP is not desirable. The harmonic mean between AP and Recall is hence adopted as a way to identify the most balanced detector. Results in Table 9, obtained with a confidence score threshold $\lambda_{od}=0.01$, show that Faster R-CNN trained on labels clustered from EgoTracks achieves the second-highest recall ($0.5224$), but also yields a very low AP value ($0.0219$). On the contrary, the YOLOv10 instance achieves the highest AP ($0.1983$) but also a low recall ($0.1634$). We attribute these skewed results to the sparse nature of annotations in EgoTracks (only a few object instances annotated per frame) and the different ways the two methods handle negative examples. The harmonic mean results of the first one are rather low ($0.0420$), while the second result is higher but not optimal at $0.1791$, highlighting the unbalanced nature of the two approaches. Training YOLOv10 on EgoObjects, which contains much denser annotations increases Recall to $0.3780$ at the expense of AP ($0.0247$), which leads to a low harmonic mean of $0.0463$. Mapping EgoTracks annotations to the EgoObjects taxonomy and training YOLOv10 using both datasets provides the most balanced detector, with a harmonic mean of $0.1882$, an AP of $0.1358$ and a Recall of $0.3067$.
YOLOWorld achieves the lowest precision value among all the methods (0.0207), but the highest recall, equal to 0.5969. Due to the low precision value, the harmonic mean has a low score equal to 0.0401. GroundingDino achieves a low AP and a reasonable recall value ($0.0450$ and $0.2572$), which is likely due to the large domain shift between data used for training GroundingDino and the challenging egocentric observations of EgoTracks. Using a Hand-Object-Detector as an object detector leads to lower results, which is because the Hand-Object-Detector only focuses on interacted objects, hence neglecting many non-interacted objects which may nevertheless be queried in the future. Lastly, RT-DETR (used in VideoAgent [15]) achieves results comparable to GroundingDino, due to similar reasons.